# Data-Centric Evolution in Autonomous Driving: A Comprehensive Survey of Big Data System, Data Mining, and Closed-Loop Technologies

### 23 Jan 2024 v1

@@ 3 Closed-Loop Data-Driven Autonomous Driving System
We’re now shifting from the previous era of software & algorithm defined autonomous driving towards the new inspiring era of big data-driven & intelligent model collaborative autonomous driving. Closed-loop data-driven systems aim to bridge the gap between AD algorithm training and their application/deployment in real-world scenarios. Unlike traditional open-loop methods, where models are passively trained on datasets collected from human client driving or road testing, closed-loop systems interact dynamically with the real environment. This approach addresses the distribution shifting challenge–where behavior learned from static datasets may not translate to the dynamic nature of real-world driving scenarios. Closed-loop systems allow AVs to learn from interactions and adapt to new situations, improving through iterative cycles of action and feedback.
However, building real-world closed-loop data-centric AD systems remain to be challenging because of several critical issues: The first issue is associated with AD data collection. In real-world data acquisition, the majority of data samples are common/normal driving scenarios, whereas corner case and abnormal driving scenario data can hardly ever be acquired. Second, further endeavors are needed to explore accurate and efficient AD data auto-labeling methods. Third, in order to alleviate the problem of AD model’s underperformance in certain scenarios within urban environments, scene data mining and scene understanding should be emphasized.
@@ 3.1 State-of-the-art Closed-Loop Autonomous Driving Pipelines
The autonomous driving industry is actively constructing integrated big data platforms in response to the challenges caused by the accumulation of massive amounts of AD data. This can be aptly termed as the new infrastructure in the era of data-driven autonomous driving. During our investigation of the data-driven closed-loop systems developed by top-tier autonomous driving companies, we have identified several commonalities. These pipelines usually follow a workflow circle that includes (1) data acquisition, (2) data storage, (3) data selection & preprocessing, (4) data labeling, (5) AD model training, (6) simulation/test validation, and (7) real-world deployment. Aside from that, the industry also emphasize the long-tail distribution problem of real-world AD datasets and the challenges when dealing with corner case.
Data, algorithms, and computing resources are the three fundamental pillars of autonomous driving technology. The quantity and quality of data determine the performance upper bound of AD algorithms/models, while computing resources serve as the vessel capacity for data processing and algorithm execution. In the realm of closed-loop data-driven autonomous driving, Tesla and NVIDIA are industry pioneers, their data system architectures offer significant reference for the development of the field. Figure 3(left) and 3 (right) give a vivid illustration of the NVIDIA MagLev system and Tesla Autopilot data close-loop platform, respectively.
@@
The NVIDIA MagLev AV platform Farabet and Koumchatzky (2020) follows the ”collect $\to$ select $\to$ label $\to$ train $\to$ test” as the programmatic, it is a reproducible workflow which enables active learning for SDC, with intelligent labeling in the loop. MagLev primarily includes two closed-loop pipelines. The first loop is autonomous driving data-centric, which starts with data ingestion and smart selection, moving through labeling and annotation, followed by model search and training. The trained models are then evaluated, debugged, and finally deployed in real-world. The second closed-loop is the platform’s infrastructure support system, which includes data center backbone and hardware infrastructure. This loop includes secure data handling, scalable DNN and system KPIs, dashboard for tracking and debugging. It supports the full cycle of AV development, ensuring continuous improvement and integration of real-world data and simulation feedback into the development process.
Tesla AutoPilot Data Platform Elluswamy (2023) is another representative AD platform which emphasize using big data-driven closed-loop pipeline to significantly boost autonomous driving model performance. The pipeline starts from source data collection, which usually comes from Tesla’s fleet learning, event-trigger vehicle-end data collection, and shadow mode. The collected data will be stored, managed, and inspected by data platform algorithms or human experts. Whenever a corner case/inaccuracy is identified, the data engine will retrieve and match data samples which are highly similar to the corner case/inaccuracy event from the existing database. At the same time, units testing will be developed to replicate the scenario and rigorously test the system’s responses. Unit tests are designed to isolate and evaluate the performance of AD model components responsible for the corner case/inaccuracy. Following that, the retrieved data samples will be labeled by either auto-labeling algorithm or human expert. Then, well-labeled data will be feedback to AD database, the database will be updated to generate new versions of training dataset for AD perception/prediction/planning/control models. After model training, validation, simulation and real-world testing, the new AD models with higher performance will be released and deployed.
@@ 3.2 High-Fidelity AD Data Generation and Simulation based on Generative AI
The majority of AD data samples collected from real-world acquisition are common/normal driving scenarios, of which we already have a massive amount of similar samples in database. However, to collect a certain type of AD data samples from real-world acquisition, we need to drive for exponentially long hours, which is not feasible in industrial applications. As such, methods for high-fidelity autonomous driving data generation and simulation have aroused significant attention from academia. CARLA Dosovitskiy et al. (2017); Deschaud (2021) is an open-source simulator for autonomous driving research that enables the generation of autonomous driving data under various settings specified by the user. The advantage of CARLA lies in its flexibility, allowing users to create diverse road conditions, traffic scenarios, and weather dynamics, which facilitates comprehensive model training and testing. However, as a simulator, its primary disadvantage lies in the domain gap. AD data generated by CARLA cannot fully mimic the real-world physics and visual effects; the dynamic and complex characteristics of real driving environment are also not represented. More recently, World Model Ha and Schmidhuber (2018); Lin et al. (2020), with its more advanced intrinsic concept and more promising performance, has been employed for high-fidelity AD data generation. World model can be defined as an AI system that builds an internal representation of the environment it perceived, and uses the learned representation to simulate data or events within the environment. The objective of general world models are to represent and simulate various situations and interactions, like a mature human will encountered in the real-world. In autonomous driving fields, GAIA-1 Hu et al. (2023) and DriveDreamer Wang et al. (2023) are representative works for data generation based on world models. GAIA-1 is a generative AI model which enables image/video to image/video generation by taking the original image/video along with text and action prompts as input. The input modalities of GAIA-1 are encoded into a unified sequence of tokens. These tokens are processed by an autoregressive transformer within a world model to predict subsequent image tokens. A video decoder then reconstructs these tokens into a coherent video output with enhanced temporal resolution, enabling dynamic and contextually rich visual content generation. DriveDreamer innovatively employs a diffusion model within its architecture, focusing on capturing the intricacies of real-world driving environments. Its two-stage training pipeline first enables the model to learn structured traffic constraints and then progresses to forecasting future states, ensuring a robust environmental understanding tailored for autonomous driving applications.
@@ 3.3 Auto-Labeling Methods for Autonomous Driving Datasets
Scene classification and understanding are foundational in autonomous driving big data platform, where the system categorizes video frames into predefined scenarios, such as driving places (streets, highway, urban overpass, main road, etc) and scene weathers (sunny, rain, snow, foggy, thunderstorm, etc). CNN-based methods are usually employed for scene classification, which include Pre-training+Fine-tune CNN models Liu et al. (2019), multi-view and multi-layer CNN models Terven et al. (2023), and various CNN-based models for improving scene representation Dixit and Vasconcelos (2016); Chen et al. (2020). Scene understanding goes beyond mere classification. It involves interpreting dynamic elements within the scene, such as surrounding vehicle agents, pedestrian, and traffic lights. Aside from image-based scene understanding, LiDAR-based data sources such as SemanticKITTI Behley et al. (2019) have also been widely adopted because of the fine-grained geometric information it provides.
@@
The emergence of 3D dynamic object auto-labeling and 3D static scene auto-labeling are to meet the the requirements of the widely adopted BEV perception technology. Waymo Qi et al. (2021) proposed a 3D Auto-Labeling pipeline from LiDAR point cloud sequential data, which employs a 3D detector to localize objects frame by frame. The identified objects’ bounding boxes across the frames are then linked through a multi-object tracker. Object track data (the corresponding point clouds + 3D bounding boxes at every frame) are extracted for each object and undergone the object-centric auto-labeling with a divide-and-conquer architecture to generate the final refined 3D bounding boxes as labels. The Auto4D pipeline Yang et al. (2021) proposed by Uber first explores AD perception labeling under spatial-temporal scale. 3D object bounding box labeling within spatial scale, along with 1D corresponded timestamp labeling within temporal scale, is termed as 4D labeling in autonomous driving field. Auto4D pipeline begins with sequential LiDAR point clouds to establish an initial object trajectory. This trajectory is refined by the Object Size Branch, which uses object observations to encode and decode object size. Concurrently, the Motion Path Branch encodes path observations and motion, allowing the Path Decoder to refine the trajectory with constant object size.
@@
3D static scene auto-labeling can be considered as HD-Map generation, where lanes, road boundary, crosswalks, traffic lights, and other related elements in the driving scene should be annotated. There are several attractive research works under this topic: vision-based methods such as MV-Map Xie et al. (2023), NeMO Zhu et al. (2023); LiDAR-based methods such as VMA Chen et al. (2023c); Pre-training 3D scene reconstruction methods such as Occ-BEV Min et al. (2023), OccNet Tong et al. (2023), AD-PT Yuan et al. (2023), ALSO Boulch et al. (2023). VMA is a recently proposed work for 3D static scene auto-labeling. VMA framework utilizes crowd-sourced, multi-trip aggregated LiDAR point clouds to reconstruct static scenes, segmenting them into units for processing. The MapTR-based unit annotator encodes the raw inputs into feature maps, through querying and decoding, yielding semantically typed point sequences. The outputs of VMA are vectorized maps, they will be refined via closed-loop annotation and human verification, resulting in satisfied HD-Maps for autonomous driving.
@@ 5 Conclusions
This survey provides the first systematic review of data-centric evolution in autonomous driving, including the big data system, data mining, and closed-loop technologies. In this survey, we start by formulating the dataset taxonomy categorized by milestone generations, reviewing the development of AD datasets throughout the historical timeline, introducing dataset acquisition, settings, and key features. Furthermore, we elaborate the closed-loop data-driven autonomous driving system from both academic and industrial perspectives. The workflow pipelines, procedures, and key technologies within data-centric closed-loop system are discussed in detail. Empirical studies are provided to demonstrate the utilization and advantages of data-centric closed-loop AD platforms for algorithm R&D and OTA upgrading. Finally, a comprehensive discussion is carried regarding the Pros and Cons of existing data-driven autonomous driving technologies, as well as future research directions, with a focus on the new datasets beyond 3rd generation, hardware supports, personalized AD recommendation, explainable autonomous driving. We also express our concerns about Generative AI models, data security and trustworthy in the future development of autonomous driving.
@@ Acknowledgments
We deeply appreciate Qiming Zhang from School of Computer Science at The University of Sydney, Guangxin Su from Data & Knowledge Research Group at University of New South Wales, Xingbo Du from ThinkLab at Shanghai Jiao Tong University for the discussion. We also sincerely thank dear colleagues from BYD Intelligent Driving R&D Centre at Dept. of IDA in Shenzhen China, especially Shitai Zhang and Liang Zhao. Lincan Li have had a good time working with you together.

### 26 Jan 2024 v2

@@ 3 Closed-Loop Data-Driven Autonomous Driving System
We’re now shifting from the previous era of software & algorithm defined autonomous driving towards the new inspiring era of big data-driven & intelligent model collaborative autonomous driving. Closed-loop data-driven systems aim to bridge the gap between AD algorithm training and their real-world application/deployment. Unlike traditional open-loop methods, where models are passively trained on datasets collected from human client driving or road testing, closed-loop systems interact dynamically with the real environment. This approach addresses the distribution shifting challenge–where behavior learned from static datasets may not translate to the dynamic nature of real-world driving scenarios. Closed-loop systems allow AVs to learn from interactions and adapt to new situations, improving through iterative cycles of action and feedback.
However, building real-world closed-loop data-centric AD systems remain challenging because of several critical issues: The first issue is associated with AD data collection. In real-world data acquisition, the majority of data samples are common/normal driving scenarios, whereas corner case and abnormal driving scenario data can hardly ever be acquired. Second, further endeavors are needed to explore accurate and efficient AD data auto-labeling methods. Third, in order to alleviate the problem of AD model’s underperformance in certain scenarios within urban environments, scene data mining and scene understanding should be emphasized.
@@ 3.1 State-of-the-art Closed-Loop Autonomous Driving Pipelines
The autonomous driving industry is actively constructing integrated big data platforms in response to the challenges caused by the accumulation of massive amounts of AD data. This can be aptly termed as the new infrastructure in the era of data-driven autonomous driving. During our investigation of the data-driven closed-loop systems developed by top-tier AD companies/research institutions, we have identified several commonalities:
  • These pipelines usually follow a workflow circle that includes: (I) data acquisition, (II) data storage, (III) data selection & preprocessing, (IV) data labeling, (V) AD model training, (VI) simulation/test validation, and (VII) real-world deployment.
  • For the design of closed-loops within the system, existing solutions either choose separately set ”Data Close-Loop” & ”Model Close-Loop”, or separately set cycles for different stages: ”Close Loop during R&D stage” and ”Close Loop during deployment stage”.
  • Aside from that, the industry also emphasizes the long-tail distribution problem of real-world AD datasets and the challenges when dealing with corner case. Tesla and NVIDIA are industry pioneers in this realm, and their data system architectures offer significant reference for the development of the field.
@@
NVIDIA MagLev AV platform Farabet and Koumchatzky (2020) (Figure 3(left)) follows the ”collect $\to$ select $\to$ label $\to$ train $\to$ test” as the programmatic, it is a reproducible workflow which enables active learning for SDC, with intelligent labeling in the loop. MagLev primarily includes two closed-loop pipelines. The first loop is autonomous driving data-centric, which starts with data ingestion and smart selection, moving through labeling and annotation, followed by model search and training. The trained models are then evaluated, debugged, and finally deployed in real-world. The second closed-loop is the platform’s infrastructure support system, which includes data center backbone and hardware infrastructure. This loop includes secure data handling, scalable DNN and system KPIs, dashboard for tracking and debugging. It supports the full cycle of AV development, ensuring continuous improvement and integration of real-world data and simulation feedback into the development process.
Tesla AutoPilot Data Platform Elluswamy (2023) (Figure 3 (right)) is another representative AD platform that emphasizes using big data-driven closed-loop pipelines to significantly boost autonomous driving model performance. The pipeline starts from source data collection, which usually comes from Tesla’s fleet learning, event-trigger vehicle-end data collection, and shadow mode. The collected data will be stored, managed, and inspected by data platform algorithms or human experts. Whenever a corner case/inaccuracy is identified, the data engine will retrieve and match data samples that are highly similar to the corner case/inaccuracy event from the existing database. At the same time, units testing will be developed to replicate the scenario and rigorously test the system’s responses. Following that, the retrieved data samples will be labeled by either auto-labeling algorithms or human experts. Then, well-labeled data will be feedback to AD database, the database will be updated to generate new versions of training dataset for AD perception/prediction/planning/control models. After model training, validation, simulation and real-world testing, the new AD models with higher performance will be released and deployed.
@@ 3.2 High-Fidelity AD Data Generation and Simulation based on Generative AI
The majority of AD data samples collected from real-world acquisition are common/normal driving scenarios, of which we already have a massive amount of similar samples in the database. However, to collect a certain type of AD data samples from real-world acquisition, we need to drive for exponentially long hours, which is not feasible in industrial applications. As such, methods for high-fidelity autonomous driving data generation and simulation have aroused significant attention from academia. CARLA Dosovitskiy et al. (2017); Deschaud (2021) is an open-source simulator for autonomous driving research that enables the generation of autonomous driving data under various settings specified by the user. The advantage of CARLA lies in its flexibility, allowing users to create diverse road conditions, traffic scenarios, and weather dynamics, which facilitates comprehensive model training and testing. However, as a simulator, its primary disadvantage lies in the domain gap. AD data generated by CARLA cannot fully mimic the real-world physics and visual effects; the dynamic and complex characteristics of real driving environment are also not represented.
More recently, World Model Ha and Schmidhuber (2018); Lin et al. (2020), with its more advanced intrinsic concept and more promising performance, has been employed for high-fidelity AD data generation. World Model can be defined as an AI system that builds an internal representation of the environment it perceives, and uses the learned representation to simulate data or events within the environment. The objective of general world models are to represent and simulate various situations and interactions, like a mature human will encountered in the real-world. In autonomous driving fields, GAIA-1 Hu et al. (2023) and DriveDreamer Wang et al. (2023) are representative works for data generation based on world models. GAIA-1 is a generative AI model that enables image/video to image/video generation by taking the original image/video along with text and action prompts as input. The input modalities of GAIA-1 are encoded into a unified sequence of tokens. These tokens are processed by an autoregressive transformer within a world model to predict subsequent image tokens. A video decoder then reconstructs these tokens into a coherent video output with enhanced temporal resolution, enabling dynamic and contextually rich visual content generation. DriveDreamer innovatively employs a diffusion model within its architecture, focusing on capturing the intricacies of real-world driving environments. Its two-stage training pipeline first enables the model to learn structured traffic constraints and then progresses to forecasting future states, ensuring a robust environmental understanding tailored for autonomous driving applications.
@@ 3.3 Auto-Labeling Methods for Autonomous Driving Datasets
High-quality data annotation is indispensable for the success and reliability of AD algorithms. Up to date, the data labeling pipelines can be characterized as three types, ranging from traditional handicraft labeling to semi-automatic labeling and then to the most advanced fully auto-labeling approach, as depicted in Figure 4. AD data labeling is usually considered as task/model specific. The workflow starts with carefully preparing the requirements of annotation task and raw dataset. Then, the next step is to generate initial annotation results using human experts, auto-labeling algorithms, or End2End large models. Following that, the annotation quality will be checked by either human experts or automatic quality-check algorithms, according to the pre-defined requirements. If the annotation results in this round fail to pass quality check, they will be sent back to the labeling loop again and repeat this annotation job, until they meet the pre-defined requirements. Finally, we can obtain the ready-to-use labeled AD dataset.
Figure 4: The detailed workflows of mainstream AD data labeling pipelines. AD data labeling is usually task/model specific, with pre-defined requirements. It’s usually not a one-time task, but a cyclical procedure.
@@
Scene classification and understanding are foundational in autonomous driving big data platform, where the system categorizes video frames into predefined scenarios, such as driving places (street, highway, urban overpass, main road, etc) and scene weathers (sunny, rain, snow, foggy, thunderstorm, etc). CNN-based methods are usually employed for scene classification, which include Pre-training+Fine-tune CNN models Liu et al. (2019), multi-view and multi-layer CNN models Terven et al. (2023), and various CNN-based models for improving scene representation Dixit and Vasconcelos (2016); Chen et al. (2020). Scene understanding Peng et al. (2023); YOLOv8 (2023) goes beyond mere classification. It involves interpreting dynamic elements within the scene, such as surrounding vehicle agents, pedestrian, and traffic lights. Aside from image-based scene understanding, LiDAR-based data sources such as SemanticKITTI Behley et al. (2019) have also been widely adopted because of the fine-grained geometric information they provide.
@@
The emergence of 3D dynamic object auto-labeling and 3D static scene auto-labeling are to meet the the requirements of the widely adopted BEV perception technology. Waymo Qi et al. (2021) proposed a 3D Auto-Labeling pipeline from LiDAR point cloud sequential data, which employs a 3D detector to localize objects frame by frame. The identified objects’ bounding boxes across the frames are then linked through a multi-object tracker. Object track data (the corresponding point clouds + 3D bounding boxes at every frame) are extracted for each object and undergone the object-centric auto-labeling with a divide-and-conquer architecture to generate the final refined 3D bounding boxes as labels. The Auto4D pipeline Yang et al. (2021) proposed by Uber first explores AD perception labeling under spatial-temporal scale. 3D object bounding box labeling within the spatial scale, along with 1D corresponded timestamp labeling within the temporal scale, is termed as 4D labeling in autonomous driving field. Auto4D pipeline begins with sequential LiDAR point clouds to establish an initial object trajectory. This trajectory is refined by the Object Size Branch, which uses object observations to encode and decode object size. Concurrently, the Motion Path Branch encodes path observations and motion, allowing the Path Decoder to refine the trajectory with constant object size.
@@
3D static scene auto-labeling can be considered as HD-Map generation, where lanes, road boundaries, crosswalks, traffic lights, and other related elements in the driving scene should be annotated. There are several attractive research works under this topic: vision-based methods such as MV-Map Xie et al. (2023), NeMO Zhu et al. (2023); LiDAR-based methods such as VMA Chen et al. (2023c); Pre-training 3D scene reconstruction methods such as Occ-BEV Min et al. (2023), OccNet Tong et al. (2023), AD-PT Yuan et al. (2023), ALSO Boulch et al. (2023). VMA is a recently proposed work for 3D static scene auto-labeling. VMA framework utilizes crowd-sourced, multi-trip aggregated LiDAR point clouds to reconstruct static scenes, segmenting them into units for processing. The MapTR-based unit annotator encodes the raw inputs into feature maps, through querying and decoding, yielding semantically typed point sequences. The outputs of VMA are vectorized maps, they will be refined via closed-loop annotation and human verification, resulting in satisfied HD-Maps for autonomous driving.
@@ 5 Conclusions
This survey provides the first systematic review of data-centric evolution in autonomous driving, including the big data system, data mining, and closed-loop technologies. In this survey, we start by formulating the dataset taxonomy categorized by milestone generations, reviewing the development of AD datasets throughout the historical timeline, introducing dataset acquisition, settings, and key features. Furthermore, we elaborate the closed-loop data-driven autonomous driving system from both academic and industrial perspectives. The workflow pipelines, procedures, and key technologies within data-centric closed-loop system are discussed in detail. Empirical studies are provided to demonstrate the utilization and advantages of data-centric closed-loop AD platforms for algorithm R&D and OTA upgrading. Finally, a comprehensive discussion is carried regarding the Pros and Cons of existing data-driven autonomous driving technologies, as well as future research directions. Emphasis is placed on the new datasets beyond 3rd generation, hardware supports, personalized AD recommendation, explainable autonomous driving. We also express our concerns about Generative AI models, data security and trustworthy in the future development of autonomous driving.
@@ Acknowledgments
We deeply appreciate Guangxin Su from Data & Knowledge Research Group at University of New South Wales, Xingbo Du from ThinkLab at Shanghai Jiao Tong University, Yinfeng Xiang from State Key Lab of Industrial Control Technology at Zhejiang University for the discussion. We also sincerely thank dear colleagues from BYD Intelligent Driving R&D Centre at Dept. of IDA in Shenzhen China, especially Shitai Zhang and Liang Zhao. Lincan Li have had a good time working with you together.